5. Use the unemployment and minimum wage data above to answer this question. a) What is the predicted unemployment rate for a state with a minimum wage of 10 dollars? How does this compare to the actual unemployment rate in Colorado? If a state increases its minimum wage from 10 dollars to 15 dollars, what is the predicted change in the unemployment rate? b) c) Compute the predicted unemployment rate and the prediction error for each state. d) Compute the R-squared for your regression line. Interpret your R-squared in a sentence.

var 101 po r The fołlowing table contains the minimum wage and unemployment rates for a sample of eight states. State 2018 Minimum | Rate (%) Wage4 Unemployment Alabama S7 Arizona Arkansas California $11 Colorado Connecticut $10 Delaware$8 $11 $8 4 S10 4 4 10 DC $13
Answer #1

5.

Wage V/S Unemployment rate y 0.3364x+0.8455 R 0.4526 4 3 E 2 0 4 10 1 1A Minimum Wage (S)

Unemployment rate = 0.3364*wage + 0.8455
A.

For a minimum wage of $10.

Predicted unemployment rate = 0.3364*10 + 0.8455

Predicted unemployment rate = 4.21%

Actual unemployment rate in Colorado = 3%

So, predicted unemployment rate in Colorado is overstated by 1.21%

B.

For a minimum wage of $15.

Predicted unemployment rate = 0.3364*15 + 0.8455 = 5.89%

Change in predicted unemployment rate = 5.89%-4.21% = 1.68%

C.

State 2018 Minimum wage ($) Unemployment rate % Predicted unemployment rate (%) Prediction error
Alabama 7 4 3.20 0.80
Arizona 11 5 4.55 0.45
Arkansas 8 3 3.54 -0.54
California 11 4 4.55 -0.55
Colorado 10 3 4.21 -1.21
Connecticut 10 4 4.21 -0.21
Delaware 8 4 3.54 0.46
DC 13 6 5.22 0.78

D.

R Square = 45.26%

The minimum wage can only explain 45.26% of the change in unemployment rate.